OpenStudy (anonymous):

Find the volume of a sphere that has a surface area of 16 sq. in. a) (32/3)(pi)cubic inches b) 8(pi)cubic inches c) (4/3)(pi)cubic inches

OpenStudy (anonymous):

Well I'll be. This question sort of has me stumped. I know the radius needs to be found, but you end up with a pi in the radius..Can't wait to see what you come up with Calcmathlete.

OpenStudy (anonymous):

Yeah, I get you. Some of these questions are typed wrong or some other type of error. Others are just a bit dumb. lol

OpenStudy (anonymous):

\[SA = 4πr^{2}\]\[V = \frac43πr^{3}\]\[16 = 4πr^{2}\]\[\frac4π = r^{2}\]\[\sqrt{\frac4π}= r\]\[\frac{2}{\sqrt{π}} = r\]\[V = \frac43πr^{3}\]\[V = \frac43π(\frac{2}{\sqrt{π}})^{3}\]\[V = \frac43π \frac{8}{π \sqrt{π}}\]\[V = \frac43\cancelπ \frac{8}{\cancelπ \sqrt{π}}\]\[V = \frac43 * \frac{8√π}{π}\]\[V = \frac{32\sqrt{π}}{3π}\]Can you solve from here as in approximate?

OpenStudy (anonymous):

Oh wow. Calc, can you remind me how that works? How does (sqrt(pi))^3 = pi*sqrt(pi) ?

OpenStudy (anonymous):

Ah..never mind. sqrt(pi)*sqrt(pi)*sqrt(pi) = pi*sqrt(pi)... Nicely done!